# ARCANE: Adaptive Routing with Caching and Aware Network Exploration

### v4

@@ D.1. ACK Coalescing Theoretical Modeling
To assess ARCANE performance under different ACK coalescing ratios, we validate it using the theoretical model from Section 5.1.
@@ D.2. Different Tiers
We aim to verify that ARCANE performs effectively with fat-tree topologies that have three tiers. This scenario poses a slightly greater challenge for ARCANE, as a single EV must manage two hops. Nonetheless, there is no intrinsic reason why ARCANE should not perform well in such a topology.
To validate this, we execute the synthetic benchmark using the symmetric topology described in Section 4.3.1, but with three tiers instead of two. The results, shown in Figure 18, indicate that ARCANE performs comparably to the two-tier topology.
Figure 18. Studying ARCANE with a 3 tiers fat tree topology.
@@ D.3. Incremental Failures
To further demonstrate the resilience of ARCANE under failures, we conducted an experiment where we incrementally failed all but one of a switch’s uplinks at 200 µs intervals. Figure 19 presents the permutation from the perspective of the failing switch, where three of the four uplinks were permanently disabled in a staggered manner. As expected, ARCANE enters freezing mode immediately after the first failure, ensuring that failing output ports are avoided. Notably, small utilization spikes are observed on the failing links when ARCANE exits freezing mode to verify if the failure has been resolved. Since the failures are permanent for the duration of the experiment, ARCANE promptly re-enters freezing mode after detecting unresolved issues.
In this drastic scenario, OPS performs 40× worse than ARCANE, primarily due to its inability to avoid broken links, resulting in numerous retransmissions and reduced congestion windows.
Figure 19. ARCANE vs. OPS in a 32 MiB permutation with incremental persistent failures.

### v5

@@ D.1. ACK Coalescing Theoretical Modeling
To assess REPS performance under different ACK coalescing ratios, we validate it using the theoretical model from Section 5.1.
@@ D.2. Different Tiers
We aim to verify that REPS performs effectively with fat-tree topologies that have three tiers. This scenario poses a slightly greater challenge for REPS, as a single EV must manage two hops. Nonetheless, there is no intrinsic reason why REPS should not perform well in such a topology.
To validate this, we execute the synthetic benchmark using the symmetric topology described in Section 4.3.1, but with three tiers instead of two. The results, shown in Figure 18, indicate that REPS performs comparably to the two-tier topology.
Figure 18. Studying REPS with a 3 tiers fat tree topology.
@@ D.3. Incremental Failures
To further demonstrate the resilience of REPS under failures, we conducted an experiment where we incrementally failed all but one of a switch’s uplinks at 200 µs intervals. Figure 19 presents the permutation from the perspective of the failing switch, where three of the four uplinks were permanently disabled in a staggered manner. As expected, REPS enters freezing mode immediately after the first failure, ensuring that failing output ports are avoided. Notably, small utilization spikes are observed on the failing links when REPS exits freezing mode to verify if the failure has been resolved. Since the failures are permanent for the duration of the experiment, REPS promptly re-enters freezing mode after detecting unresolved issues.
In this drastic scenario, OPS performs 40× worse than REPS, primarily due to its inability to avoid broken links, resulting in numerous retransmissions and reduced congestion windows.
Figure 19. REPS vs. OPS in a 32 MiB permutation with incremental persistent failures.